Consider two production lines with fraction of nonconformities in accordance with LSL, USL.

Production Line 1

P1=0.2

Production Line 2

P2=0.2

  1. If we have a lot which is made up of 100 products (50 products from each production line), what is P(x<=1), where x is number of nonconforming product in a sample size of 10 products? (No approximation methods, Calculation with three digits after decimal point)?.
Answer #1

Here X ~B(n=10, P=0.2)

P(X1) = P(X=0) +P(X=1)

=BINOMDIST(1,10,0.2,1)

This function can be used in excel to get required answer.